Timothy's age in 13 yers will be twice what it was 15 years ago. How old is Timothy now?
Answer by raisetoana(4)

Let T be Timothy's age now.
T + 13 is Timothy's Age 13 years from now.
T - 15 is Timothy's Age 15 years ago.
T + 13 = 2(T-15)
T + 13 = 2T - 30
43 = T
Timothy is 43 years old.